OpenStudy (anonymous):

Let f(x) = 1/x and g(x) = x^2 - 9x. What two numbers are not in the domain of f^og? Separate your answers with a comma.

OpenStudy (anonymous):

Can someone help me with this?

OpenStudy (welshfella):

Hint a denominator cannot be zero

OpenStudy (welshfella):

f^og = 1 / (x^2 - 9x)

OpenStudy (welshfella):

so you need to find values of x which will make x ^2 - 9x = 0

OpenStudy (welshfella):

solve x^2 - 9x = 0

OpenStudy (anonymous):

Sorry I can't figure this out

OpenStudy (welshfella):

can you factor x^2 - 9x what is common to x^2 and 9x ?

OpenStudy (welshfella):

what is in x^2 and 9 x ?

OpenStudy (anonymous):

....x?